# Superconducting pairing correlations on a trapped-ion quantum computer

### v2

@@ Abstract
The Fermi-Hubbard model is the starting point for the simulation of many strongly correlated materials, including high-temperature superconductors, whose modelling is a key motivation for the construction of quantum simulation and computing devices [arovas_hubbard_2022, bloch_quantum_2012]. However, the detection of superconducting pairing correlations has so far remained out of reach, both because of their off-diagonal character—which makes them inaccessible to local density measurements—and because of the difficulty of preparing superconducting states [schlomer_local_2024]. Here, we report measurement of significant pairing correlations in three different regimes of Fermi-Hubbard models simulated on Quantinuum’s Helios trapped-ion quantum computer. Specifically, we measure non-equilibrium pairing induced by an electromagnetic field in the half-filled square lattice model, $d$-wave pairing in an approximate ground state of the checkerboard Hubbard model at $1/6$-doping, and $s$-wave pairing in a bilayer model relevant to nickelate superconductors. These results show that a quantum computer can reliably create and probe physically relevant states with superconducting pairing correlations, opening a path to the exploration of superconductivity with quantum computers.
@@ I Introduction
Despite decades of study and increasing technological relevance [coombs_high-temperature_2024], the mechanism underlying unconventional superconductivity remains unresolved. Theoretical modelling can be invaluable in understanding these materials and informing experimental and technological progress. The starting point for modelling the electronic correlations in many of these systems is the Fermi-Hubbard model
@@
where $t_{ij}$ describes spinful fermions hopping between proximate orbitals on a lattice and $U$ models an on-site Coulomb repulsion. The ability to compute arbitrary observables in this model, including superconducting pairing correlations, would be highly valuable for elucidating the basic mechanisms of superconductivity and connecting with experiments [qin2022hubbard].
In the pursuit of this goal, remarkable progress has been made in the development of classical computer simulations, to the point where different methods are starting to agree on ground state predictions for certain experimentally relevant parameters [xu_coexistence_2024, simons_collaboration_on_the_many-electron_problem_absence_2020]. Yet, extending these capabilities to simulate thermal and non-equilibrium properties essential to phenomena like strange metallicity [bruin_linear_resistivity] and light-induced superconductivity [fava_magnetic_2024] remains a formidable challenge. Indeed, the difficulty of classically simulating quantum many-body systems of this kind was one of the initial motivations to develop quantum simulation platforms, which promise to capture their dynamics in an unbiased way using only polynomially scaling resources. Such devices have become experimental reality, with purpose-built analog simulators now challenging classical simulation techniques in low-but-finite temperature regimes of the Hubbard model [xu_neutral-atom_2025, bourgund_formation_2025, kendrick2025pseudogapfermihubbardquantumsimulator, chalopin2024probingmagneticoriginpseudogap], and digital quantum computers making fast progress on their path towards fault-tolerance [dasu_breaking_2025, lacroix_scaling_2025, sales_rodriguez_experimental_2025]. In the context of superconducitivity, using these platforms to produce and measure a state with superconducting pairing correlations is an essential step towards realising their potential. However, despite the existence of experimental proposals [mark2025efficiently, schlomer_local_2024], to the best of our knowledge, no quantum simulation platform has observed superconducting pairing correlations in the repulsive Fermi-Hubbard model to date.
@@
Here, we report measurement of non-zero pairing correlations in three regimes of the Hubbard model, simulated on a trapped-ion quantum computer. First, we show that electromagnetic radiation can induce non-equilibrium $\eta$-type [yang_ensurematheta_1989] pairing correlations in a half-filled square lattice, whose ground-state is a non-superconducting antiferromagnet. Second, we prepare an approximate ground state of the $1/6$-doped checkerboard model in which $2\times 2$ Hubbard plaquettes are weakly coupled [yao_myriad_2007] and observe singlet pairing correlations with $d$-wave symmetry. Finally, we prepare low-energy states of a bilayer Hubbard model relevant to the recently discovered nickelate superconductor $\mathrm{La}_{3}\mathrm{Ni}_{2}\mathrm{O}$ [sun_signatures_2023], and observe $s$-wave pairing in the limit of strong interlayer spin-exchange coupling.
All of our experiments are carried out on Quantinuum’s Helios quantum computer [to_be_published_helios_nodate], which operates on 98 effectively all-to-all connected $^{137}\mathrm{Ba}^{+}$ hyperfine qubits. Our results are enabled by several theoretical advancements: (i) we develop ground state preparation techniques based on the injection of states with locally fixed fermionic parity, (ii) we use new time evolution circuits for hopping and exchange terms that balance Trotter error and gate count, and (iii) we adapt and extend local fermionic encodings  [derby_compact_2021, derby2021compactfermionqubitmapping] and develop new error mitigation strategies.
@@ II Half-Filling
To benchmark the performance of quantum hardware and software, we begin our experiments with the $6\times 6$ square lattice Hubbard model at half-filling and nearest-neighbour hopping, which offers opportunities to compare with exact classical simulations. To this end, we first set $t=1$ and $U=0$, initialise a state with 36 non-interacting fermions in a region $\mathcal{A}$ and track the population imbalance between $\mathcal{A}$ and its complement $\overline{\mathcal{A}}$, denoted $I_{\mathcal{A}}=n_{\mathcal{A}}-n_{\overline{\mathcal{A}}}$, over time. To carry out the evolution we use a Trotter circuit $U_{\mathrm{hop}}(\tau t)$ with $\tau$ some Trotter step, which approximates $e^{i\tau t\sum_{\langle ij\rangle\sigma}c^{\dagger}_{i\sigma}c_{j\sigma}+\mathrm{h.c.}}$ (see Section  S4). Here, and in all other experiments, we use mixed periodic boundary conditions (mPBC), i.e. we average our results over periodic and anti-periodic boundary conditions in the $x$- and $y$-direction, which minimises finite-size effects [lin_twist-averaged_2001] (see Section S13).
In order to capture the fermionic statistics, we use a local fermionic encoding [derby_compact_2021, derby2021compactfermionqubitmapping, nigmatullin_experimental_2025], which requires us to prepare a toric code state on 18 ancillae prior to evolution (see Sections S2 and S5). The 18 stabilisers of the toric code commute with all fermionic bilinears and we choose to measure them simultaneously with $I_{\mathcal{A}}$. As such, each shot comes with diagnostic information about how much noise has occurred. A simple strategy to exploit that information is to group shots into two sets: those with lower-than-median number of violated stabilisers (“good” shots) and the complement (“bad” shots). Extrapolating the corresponding expectation values yields an error mitigation strategy that avoids the need to run additional, noise-amplified circuits (see Section S3 for more details).
@@
Figure 2: $D$-wave pairing in the doped checkerboard model. (a) Using a brickwall circuit, an approximate ground state of the ferromagnetic $3\times 3$ lattice XXZ model is prepared. (b) The state is injected into the fermionic Fock space via a local isometry $V_{\mathrm{plaquette}}$ which maps the qubit states to the ground states of the $2\times 2$ Hubbard model in the sector with two $(s)$ and four $(d)$ fermions, respectively. (c) The resulting wave function is a superposition of states with three $s$- and six $d$-plaquettes (one configuration shown). It approximates the ground state of the $6\times 6$ square lattice with inhomogeneous hopping $t\gg t^{\prime}$, $U/t=2$ and $1/6$ doping. (d) Pairing correlations (8) of the approximate ground state which show $d$-wave symmetry. For bonds on the left of a strong plaquette $P_{b}(x,y)=P_{b}(-x,-y)$ by definition, but all bonds are shown for visual completeness. Avg. (max.) standard error on the mean is 0.016 (0.017). (e) Sketch of the phase diagram of the checkerboard model at low doping, as conjectured in [tsai_optimal_2008], with the prepared state indicated by the red circle (left). We evolve the prepared state by a one-step adiabatic evolution towards U/t=8,t/t=1/2U/t=8,\ $^{\prime}/t$=1/2 and measure the energy density with respect to the target parameters (right).
After using the exactly solvable limit for validation, we now turn to the interacting case. Specifically, we aim to prepare a low-energy state in the half-filled model at $U/t=8$, which is experimentally relevant to cuprate materials and a standard benchmark for numerical simulations and analog simulator platforms [bohrdt2021exploration, simons_collaboration_on_the_many-electron_problem_solutions_2015]. To this end, we approach the model from its $J=4t^{2}/U\rightarrow 0$ limit, at which point the low-energy physics is captured by the isotropic ($\delta=1$) Heisenberg limit of the XXZ model
@@
The results of this strategy are shown in Fig. 1d. The lowest raw (mitigated) energy density is $\langle H\rangle/N-U/4=-2.36\pm 0.02$ $(-2.43\pm 0.04)$ and is achieved for $(D_{\mathrm{Heisenberg}},D_{\mathrm{Hubbard}})=(1,1)$, indicating that increasing noise may overcompensate for the increase in variational power of deeper circuits. The ground state energy for periodic boundary conditions is approximately 2.5278-2.5278 [qin_qmc_benchmark_2016]. Note that, throughout this work, we set the energy scale such that the fully depolarised state ($\rho\propto\mathds{1}$) has energy zero.
@@
Nevertheless, these low-energy states can be used as a resource to investigate non-equilibrium phenomena, including light-induced superconductivity. Specifically, Ref. [kaneko_photoinduced_2019] showed that the antiferromagnetic ground states of a 14-site Fermi-Hubbard chain develop doublon pairing correlations
@@
when exposed to radiation in a range of frequencies around $\sim U=8t$, where $\Delta^{\dagger}_{i}=c^{\dagger}_{i\uparrow}c^{\dagger}_{i\downarrow}$ [yang_ensurematheta_1989]. Moreover, the sign of $P_{\eta}$ was found to follow a staggered pattern which stabilises when turning off the radiation and even increases for larger distances.
@@
Fig. 1g shows a significant increase of overall pairing correlations induced by the pulse. These correlations are especially strong for nearest neighbours and occur predominantly in a staggered pattern. The staggered average $P^{\mathrm{stag}}_{\eta}=\sum_{x,y\neq 0,0}(-1)^{\mathrm{sublattice}}P_{\eta}(x,y)$ decreases during a subsequent relaxation implemented via two more Trotter steps of size $\tau=0.375$ with the field turned off, although it remains significantly larger than the pre-pulse value. Contributions to the staggered average from nearest neighbours decrease during the relaxation, while those from larger distances increase (see Section S11). These results provide evidence that the phenomena observed in Ref. [kaneko_photoinduced_2019] can carry over to larger lattices, and may even be robust to initial states with non-zero energy density above the ground state and to Hamiltonian imperfections caused by finite Trotter step size.
@@ III The doped checkerboard model
While we have found pairing correlations in a half-filled model in the previous section, modelling light-matter interactions in cuprates almost certainly requires accounting for lattice distortions as well as dopants [tsai_superconductivity_2006, zhang_photoinduced_2023, fava_magnetic_2024]. To study both of these concepts, we now introduce 1/6 hole dopants into the $6\times 6$ checkerboard model [tsai_superconductivity_2006, yao_myriad_2007], in which the lattice is partitioned into non-overlapping $2\times 2$ plaquettes and $t_{ij}=t$ ($t^{\prime}$) if $\langle ij\rangle$ is a strong (weak) bond within (between) the plaquettes (Fig. 2c). Lattice inhomogeneities of this kind can either be a natural part of the material structure or exist transiently due to lattice vibrations. An exact diagonalisation study on a $4\times 4$-lattice showed the checkerboard model to potentially host a $d$-wave superconducting phase in a large region of parameter space (Fig. 2e) [tsai_optimal_2008].
To access this phase experimentally, we follow a strategy inspired by Ref. [rey_controlled_2009]. Specifically, we consider the limit $t^{\prime}=0,U/t=2$ at which the plaquettes decouple. The two lowest-energy states of the plaquette are given by a state $|s\rangle$ with $s$-wave symmetry in the quarter-filled sector and a state $|d\rangle$ with $d$-wave symmetry in the half-filled sector [scalapino_local_1996]. At 1/6-doping on the $6\times 6$ lattice, all states with three $|s\rangle$ and six $|d\rangle$ plaquettes are degenerate ground states in this limit. The degeneracy is lifted when turning on hopping $t^{\prime}>0$ between the plaquettes. The unique perturbative ground state that emerges is precisely the ground state of the ferromagnetic XXZ model $H_{\mathrm{XXZ}}$ at $\delta\approx-1$ and $\sum_{i}Z_{i}=-3$ on a $3\times 3$ lattice (see Section S6).
@@ IV A Bilayer Hubbard Model
One advantage of an effectively all-to-all connected quantum computer is the ability to simulate arbitrary geometries without changes to the experimental setup. We utilize this capability to simulate a model of the recently discovered bilayer nickelate superconductor $\mathrm{La}_{3}\mathrm{Ni}_{2}\mathrm{O}$ [sun_signatures_2023]. This material has been modelled by two quarter-filled square lattice Hubbard models (1) with nearest-neighbour hopping and strong inter-layer exchange coupling
@@
where $H_{\mathrm{exchange}}=J\sum_{i}\mathbf{S}_{iA}\cdot\mathbf{S}_{iB}-n_{iA}n_{iB}/4$ and $A$ and $B$ denote the two layers of the system [lu_interlayer-coupling-driven_2024, qu_bilayer_2024, schlomer_superconductivity_2024, oh_type-ii_2023]. In the limit $t/J\rightarrow 0$ (and $U\geq 0$), the ground state subspace is spanned by tightly bound singlets between the layers (Fig. 3c). Second-order degenerate perturbation theory again leads to a single-layer effective ferromagnetic XXZ model (2), in this instance with anisotropy $\delta=-2/3$.
@@
In reality, the relevant parameters describing $\mathrm{La}_{3}\mathrm{Ni}_{2}\mathrm{O}$ deviate from the perturbative limit, likely being closer to $t/J=0.7,U/t=10$ [lu_interlayer-coupling-driven_2024, qu_bilayer_2024] and the tightly coupled singlet pairs were proposed to undergo a BEC-BCS crossover at $t/J\sim 0.6$ in the related $t_{\parallel}-J_{\perp}-J_{\parallel}$ model [schlomer_superconductivity_2024]. In our final set of experiments, we study the feasibility of an adiabatic evolution from the perturbative ground state towards those parameters. To that end, we consider a short linear ramp of the parameters $(t/J,U/J)=(0,0)\rightarrow(0.7,7)$ while keeping $J=1$, simulated via $M=1,2$ Trotter steps
@@
where $\tau=0.4$. The relatively large Trotter step size is afforded by the use of novel circuit gadgets (shown in Section S4) to implement the exchange coupling $U_{\mathrm{exchange}}(\tau J)=e^{-i\tau J\sum_{i}\mathbf{S}_{iA}\cdot\mathbf{S}_{iB}-n_{iA}n_{iB}/4}$ directly, rather than via a high-frequency expansion [schlomer_local_2024, bourgund_formation_2025]. Fig. 3e shows that, although the $M=1$ strategy succeeds at lowering the energy with respect to the target Hamiltonian at $t/J=0.7$, $U/t=10$, heating from hardware errors eventually outcompetes adiabatic cooling for longer evolution times. Nearest-neighbour pairing correlations involving rungs orthogonal to the $A$- and $B$- layers decrease rapidly during the adiabatic evolution, although correlations between “slanted” singlets that involve coordinates $(x,y)$ in the $A$ layer and $(x+1,y)$ in the $B$ layer are likely to increase, indicating that the coupling between the pairs becomes less tight.

### v3

@@ Abstract
The Fermi-Hubbard model is the starting point for the simulation of many strongly correlated materials, including high-temperature superconductors, whose modelling is a key motivation for the construction of quantum simulation and computing devices [6, 13]. However, the detection of superconducting pairing correlations has so far remained out of reach, both because of their off-diagonal character—which makes them inaccessible to local density measurements—and because of the difficulty of preparing superconducting states [71]. Here, we report measurement of significant pairing correlations in three different regimes of Fermi-Hubbard models simulated on Quantinuum’s Helios trapped-ion quantum computer. Specifically, we measure non-equilibrium pairing induced by an electromagnetic field in the half-filled square lattice model, $d$-wave pairing in an approximate ground state of the checkerboard Hubbard model at $1/6$-doping, and $s$-wave pairing in a bilayer model relevant to nickelate superconductors. These results show that a quantum computer can reliably create and probe physically relevant states with superconducting pairing correlations, opening a path to the exploration of superconductivity with quantum computers.
@@ I Introduction
Despite decades of study and increasing technological relevance [22], the mechanism underlying unconventional superconductivity remains unresolved. Theoretical modelling can be invaluable in understanding these materials and informing experimental and technological progress. The starting point for modelling the electronic correlations in many of these systems is the Fermi-Hubbard model
@@
where $t_{ij}$ describes spinful fermions hopping between proximate orbitals on a lattice and $U$ models an on-site Coulomb repulsion. The ability to compute arbitrary observables in this model, including superconducting pairing correlations, would be highly valuable for elucidating the basic mechanisms of superconductivity and connecting with experiments [63].
In the pursuit of this goal, remarkable progress has been made in the development of classical computer simulations, to the point where different methods are starting to agree on ground state predictions for certain experimentally relevant parameters [83, 62]. Yet, extending these capabilities to simulate thermal and non-equilibrium properties essential to phenomena like strange metallicity [17] and light-induced superconductivity [31] remains a formidable challenge. Indeed, the difficulty of classically simulating quantum many-body systems of this kind was one of the initial motivations to develop quantum simulation platforms, which promise to capture their dynamics in an unbiased way using only polynomially scaling resources. Such devices have become experimental reality, with purpose-built analog simulators now challenging classical simulation techniques in low-but-finite temperature regimes of the Hubbard model [84, 15, 42, 18], and digital quantum computers making fast progress on their path towards fault-tolerance [23, 45, 69]. In the context of superconducitivity, using these platforms to produce and measure a state with superconducting pairing correlations is an essential step towards realising their potential. However, despite the existence of experimental proposals [53, 71], to the best of our knowledge, no quantum simulation platform has observed superconducting pairing correlations in the repulsive Fermi-Hubbard model to date.
@@
Here, we report measurement of non-zero pairing correlations in three regimes of the Hubbard model, simulated on a trapped-ion quantum computer. First, we show that electromagnetic radiation can induce non-equilibrium $\eta$-type [85] pairing correlations in a half-filled square lattice, whose ground-state is a non-superconducting antiferromagnet. Second, we prepare an approximate ground state of the $1/6$-doped checkerboard model in which $2\times 2$ Hubbard plaquettes are weakly coupled [86] and observe singlet pairing correlations with $d$-wave symmetry. Finally, we prepare low-energy states of a bilayer Hubbard model relevant to the recently discovered nickelate superconductor $\mathrm{La}_{3}\mathrm{Ni}_{2}\mathrm{O}$ [75], and observe $s$-wave pairing in the limit of strong interlayer spin-exchange coupling.
All of our experiments are carried out on Quantinuum’s Helios quantum computer [5], which operates on 98 effectively all-to-all connected $^{137}\mathrm{Ba}^{+}$ hyperfine qubits. Our results are enabled by several theoretical advancements: (i) we develop ground state preparation techniques based on the injection of states with locally fixed fermionic parity, (ii) we use new time evolution circuits for hopping and exchange terms that balance Trotter error and gate count, and (iii) we adapt and extend local fermionic encodings  [24, 25] and develop new error mitigation strategies.
@@ II Half-Filling
To benchmark the performance of quantum hardware and software, we begin our experiments with the $6\times 6$ square lattice Hubbard model at half-filling and nearest-neighbour hopping, which offers opportunities to compare with exact classical simulations. To this end, we first set $t=1$ and $U=0$, initialise a state with 36 non-interacting fermions in a region $\mathcal{A}$ and track the population imbalance between $\mathcal{A}$ and its complement $\overline{\mathcal{A}}$, denoted $I_{\mathcal{A}}=n_{\mathcal{A}}-n_{\overline{\mathcal{A}}}$, over time. To carry out the evolution we use a Trotter circuit $U_{\mathrm{hop}}(\tau t)$ with $\tau$ some Trotter step, which approximates $e^{i\tau t\sum_{\langle ij\rangle\sigma}c^{\dagger}_{i\sigma}c_{j\sigma}+\mathrm{h.c.}}$ (see Section  S4). Here, and in all other experiments, we use mixed periodic boundary conditions (mPBC), i.e. we average our results over periodic and anti-periodic boundary conditions in the $x$- and $y$-direction, which minimises finite-size effects [49] (see Section S13).
In order to capture the fermionic statistics, we use a local fermionic encoding [24, 25, 59], which requires us to prepare a toric code state on 18 ancillae prior to evolution (see Sections S2 and S5). The 18 stabilisers of the toric code commute with all fermionic bilinears and we choose to measure them simultaneously with $I_{\mathcal{A}}$. As such, each shot comes with diagnostic information about how much noise has occurred. A simple strategy to exploit that information is to group shots into two sets: those with lower-than-median number of violated stabilisers (“good” shots) and the complement (“bad” shots). Extrapolating the corresponding expectation values yields an error mitigation strategy that avoids the need to run additional, noise-amplified circuits (see Section S3 for more details).
@@
Figure 2: $D$-wave pairing in the doped checkerboard model. (a) Using a brickwall circuit, an approximate ground state of the ferromagnetic $3\times 3$ lattice XXZ model is prepared. (b) The state is injected into the fermionic Fock space via a local isometry $V_{\mathrm{plaquette}}$ which maps the qubit states to the ground states of the $2\times 2$ Hubbard model in the sector with two $(s)$ and four $(d)$ fermions, respectively. (c) The resulting wave function is a superposition of states with three $s$- and six $d$-plaquettes (one configuration shown). It approximates the ground state of the $6\times 6$ square lattice with inhomogeneous hopping $t\gg t^{\prime}$, $U/t=2$ and $1/6$ doping. (d) Pairing correlations (8) of the approximate ground state which show $d$-wave symmetry. For bonds on the left of a strong plaquette $P_{b}(x,y)=P_{b}(-x,-y)$ by definition, but all bonds are shown for visual completeness. Avg. (max.) standard error on the mean is 0.016 (0.017). (e) Sketch of the phase diagram of the checkerboard model at low doping, as conjectured in [78], with the prepared state indicated by the red circle (left). We evolve the prepared state by a one-step adiabatic evolution towards U/t=8,t/t=1/2U/t=8,\ $^{\prime}/t$=1/2 and measure the energy density with respect to the target parameters (right).
After using the exactly solvable limit for validation, we now turn to the interacting case. Specifically, we aim to prepare a low-energy state in the half-filled model at $U/t=8$, which is experimentally relevant to cuprate materials and a standard benchmark for numerical simulations and analog simulator platforms [14, 46]. To this end, we approach the model from its $J=4t^{2}/U\rightarrow 0$ limit, at which point the low-energy physics is captured by the isotropic ($\delta=1$) Heisenberg limit of the XXZ model
@@
The results of this strategy are shown in Fig. 1d. The lowest raw (mitigated) energy density is $\langle H\rangle/N-U/4=-2.36\pm 0.02$ $(-2.43\pm 0.04)$ and is achieved for $(D_{\mathrm{Heisenberg}},D_{\mathrm{Hubbard}})=(1,1)$, indicating that increasing noise may overcompensate for the increase in variational power of deeper circuits. The ground state energy for periodic boundary conditions is approximately 2.5278-2.5278 [64]. Note that, throughout this work, we set the energy scale such that the fully depolarised state ($\rho\propto\mathds{1}$) has energy zero.
@@
Nevertheless, these low-energy states can be used as a resource to investigate non-equilibrium phenomena, including light-induced superconductivity. Specifically, Ref. [41] showed that the antiferromagnetic ground states of a 14-site Fermi-Hubbard chain develop doublon pairing correlations
@@
when exposed to radiation in a range of frequencies around $\sim U=8t$, where $\Delta^{\dagger}_{i}=c^{\dagger}_{i\uparrow}c^{\dagger}_{i\downarrow}$ [85]. Moreover, the sign of $P_{\eta}$ was found to follow a staggered pattern which stabilises when turning off the radiation and even increases for larger distances.
@@
Fig. 1g shows a significant increase of overall pairing correlations induced by the pulse. These correlations are especially strong for nearest neighbours and occur predominantly in a staggered pattern. The staggered average $P^{\mathrm{stag}}_{\eta}=\sum_{x,y\neq 0,0}(-1)^{\mathrm{sublattice}}P_{\eta}(x,y)$ decreases during a subsequent relaxation implemented via two more Trotter steps of size $\tau=0.375$ with the field turned off, although it remains significantly larger than the pre-pulse value. Contributions to the staggered average from nearest neighbours decrease during the relaxation, while those from larger distances increase (see Section S11). These results provide evidence that the phenomena observed in Ref. [41] can carry over to larger lattices, and may even be robust to initial states with non-zero energy density above the ground state and to Hamiltonian imperfections caused by finite Trotter step size.
@@ III The doped checkerboard model
While we have found pairing correlations in a half-filled model in the previous section, modelling light-matter interactions in cuprates almost certainly requires accounting for lattice distortions as well as dopants [77, 87, 31]. To study both of these concepts, we now introduce 1/6 hole dopants into the $6\times 6$ checkerboard model [77, 86], in which the lattice is partitioned into non-overlapping $2\times 2$ plaquettes and $t_{ij}=t$ ($t^{\prime}$) if $\langle ij\rangle$ is a strong (weak) bond within (between) the plaquettes (Fig. 2c). Lattice inhomogeneities of this kind can either be a natural part of the material structure or exist transiently due to lattice vibrations. An exact diagonalisation study on a $4\times 4$-lattice showed the checkerboard model to potentially host a $d$-wave superconducting phase in a large region of parameter space (Fig. 2e) [78].
To access this phase experimentally, we follow a strategy inspired by Ref. [66]. Specifically, we consider the limit $t^{\prime}=0,U/t=2$ at which the plaquettes decouple. The two lowest-energy states of the plaquette are given by a state $|s\rangle$ with $s$-wave symmetry in the quarter-filled sector and a state $|d\rangle$ with $d$-wave symmetry in the half-filled sector [70]. At 1/6-doping on the $6\times 6$ lattice, all states with three $|s\rangle$ and six $|d\rangle$ plaquettes are degenerate ground states in this limit. The degeneracy is lifted when turning on hopping $t^{\prime}>0$ between the plaquettes. The unique perturbative ground state that emerges is precisely the ground state of the ferromagnetic XXZ model $H_{\mathrm{XXZ}}$ at $\delta\approx-1$ and $\sum_{i}Z_{i}=-3$ on a $3\times 3$ lattice (see Section S6).
@@ IV A Bilayer Hubbard Model
One advantage of an effectively all-to-all connected quantum computer is the ability to simulate arbitrary geometries without changes to the experimental setup. We utilize this capability to simulate a model of the recently discovered bilayer nickelate superconductor $\mathrm{La}_{3}\mathrm{Ni}_{2}\mathrm{O}$ [75]. This material has been modelled by two quarter-filled square lattice Hubbard models (1) with nearest-neighbour hopping and strong inter-layer exchange coupling
@@
where $H_{\mathrm{exchange}}=J\sum_{i}\mathbf{S}_{iA}\cdot\mathbf{S}_{iB}-n_{iA}n_{iB}/4$ and $A$ and $B$ denote the two layers of the system [52, 65, 72, 60]. In the limit $t/J\rightarrow 0$ (and $U\geq 0$), the ground state subspace is spanned by tightly bound singlets between the layers (Fig. 3c). Second-order degenerate perturbation theory again leads to a single-layer effective ferromagnetic XXZ model (2), in this instance with anisotropy $\delta=-2/3$.
@@
In reality, the relevant parameters describing $\mathrm{La}_{3}\mathrm{Ni}_{2}\mathrm{O}$ deviate from the perturbative limit, likely being closer to $t/J=0.7,U/t=10$ [52, 65] and the tightly coupled singlet pairs were proposed to undergo a BEC-BCS crossover at $t/J\sim 0.6$ in the related $t_{\parallel}-J_{\perp}-J_{\parallel}$ model [72]. In our final set of experiments, we study the feasibility of an adiabatic evolution from the perturbative ground state towards those parameters. To that end, we consider a short linear ramp of the parameters $(t/J,U/J)=(0,0)\rightarrow(0.7,7)$ while keeping $J=1$, simulated via $M=1,2$ Trotter steps
@@
where $\tau=0.4$. The relatively large Trotter step size is afforded by the use of novel circuit gadgets (shown in Section S4) to implement the exchange coupling $U_{\mathrm{exchange}}(\tau J)=e^{-i\tau J\sum_{i}\mathbf{S}_{iA}\cdot\mathbf{S}_{iB}-n_{iA}n_{iB}/4}$ directly, rather than via a high-frequency expansion [71, 15]. Fig. 3e shows that, although the $M=1$ strategy succeeds at lowering the energy with respect to the target Hamiltonian at $t/J=0.7$, $U/t=10$, heating from hardware errors eventually outcompetes adiabatic cooling for longer evolution times. Nearest-neighbour pairing correlations involving rungs orthogonal to the $A$- and $B$- layers decrease rapidly during the adiabatic evolution, although correlations between “slanted” singlets that involve coordinates $(x,y)$ in the $A$ layer and $(x+1,y)$ in the $B$ layer are likely to increase, indicating that the coupling between the pairs becomes less tight.